Cheryl and Judy trudged at 4 mph until their packs got too heavy. Then they dropped their packs and continued at the brisk pace of 6 mph. If the total trip of 56 miles took 12 hours, how long did they trudge? How long did they walk briskly.
Answer by stanbon(75887)

Cheryl and Judy trudged at 4 mph until their packs got too heavy. Then they dropped their packs and continued at the brisk pace of 6 mph. If the total trip of 56 miles took 12 hours, how long did they trudge? How long did they walk briskly.
-----
Trudge DATA:
rate = 4 mph ; distance = x miles ; time = x/4 hrs
----------------------
Brisk Pace DATA:
rate = 6 mph ; distance = 56-x miles ; time = (56-x)/6 hrs
---------
Equation:
time + time = 12 hrs.
x/4 + (56-x)/6 = 12
6x +4(56-x) = 24*12
----
6x + 224 - 4x = 24*12
2x = 64
x = 32 miles (trudge distance)
56-x = 24 miles (brisk pace distance)
==========================================
Cheers,
Stan H.
====================